An electron is placed in an electric potential of 1nV. Determine the velocity of the electron after passing through the potential assuming it was motionless in the beginning

1
Expert's answer
2022-03-22T15:47:12-0400

We know that

"\frac{1}{2}mv^2=eV"

"v=\sqrt{\frac{2eV}{m}}"


"v=\sqrt{\frac{2\times1.6\times10^{-19}\times10^{-9}}{9.1\times10^{-31}}}=18.75m/sec"